# CleanVul: Automatic Function-Level Vulnerability Detection in Code Commits Using LLM Heuristics

### v3

@@ Abstract.
To address these challenges, this paper introduces the first methodology that leverages the Large Language Model (LLM) with a heuristic enhancement to automatically identify vulnerability-fixing changes from VFCs, achieving an F1-score of 0.82. VulSifter was applied to a large-scale study, where we conducted a comprehensive crawl of 127,063 repositories on GitHub, resulting in the acquisition of 5,352,105 commits. The LLM with heuristic enhancement approach involves utilizing an LLM to comprehend code semantics and contextual information, while applying heuristics to filter out unrelated changes. We developed CleanVul, a high-quality dataset comprising 11,632 functions using our LLM heuristic enhancement approach, demonstrating Correctness (90.6%) comparable to established datasets such as SVEN (94.0%) (He and Vechev, 2023) and PrimeVul (86.0%) (Ding et al., 2024). In this context, Correctness is defined as the percentage of genuine vulnerable functions in the vulnerability dataset, which helps assess the effectiveness of VulSifter in identifying and filtering vulnerability-fixing changes. To evaluate the effectiveness of our CleanVul dataset, we conducted experiments focusing on fine-tuning various LLMs on the CleanVul dataset and other high-quality datasets. Our evaluation was conducted on multiple popular programming languages, including Java, Python, JavaScript, C#, C, and C++. Evaluation results reveal that LLMs fine-tuned on CleanVul not only exhibit enhanced accuracy but also superior generalization capabilities compared to those trained on uncleaned datasets. Specifically, models trained on CleanVul and tested on PrimeVul achieve accuracy higher than those trained and tested exclusively on PrimeVul, validating the effectiveness of our methodology.
@@ 1. Introduction
This paper introduces the first methodology to address the gap in accurately identifying vulnerability-related changes within VFCs. We first examine the reasons why many changes cataloged as vulnerability fixes do not pertain to actual security vulnerabilities. Our empirical study reveals that approximately 80% of non-vulnerability changes consist of test-related modifications and general bug fixes, with additional changes spanning support updates, code refactoring, and documentation updates. Building on this analysis, we developed VulSifter, a novel approach that combines LLM with heuristic filtering to automatically analyze VFCs and eliminate noise, resulting in a refined vulnerability dataset called CleanVul. By applying VulSifter, the Correctness of genuine vulnerability fixes in the dataset improved from 28.7% to 90.6%. Correctness is defined as the percentage of genuine vulnerable functions in the vulnerability dataset, and it helps assess the effectiveness of VulSifter in identifying and filtering vulnerability-fixing changes. CleanVul comprises 11,632 functions, both vulnerable and benign, and achieves a level of Correctness (90.6%) comparable to established datasets such as SVEN (He and Vechev, 2023) and PrimeVul (Ding et al., 2024). For comparison, the manually curated SVEN dataset contains 803 functions with 94.0% Correctness, while PrimeVul, which identifies vulnerabilities by matching function names with NVD descriptions, contains 6,968 NVD-sourced functions with 86.0% Correctness. As the first automated approach to improve dataset Correctness without requiring NVD linkage or other constraints, CleanVul represents a significant advancement in function-level vulnerability detection. It complements PrimeVul’s NVD-based approach by automatically filtering noise from GitHub VFCs to obtain high-quality vulnerability datasets.
@@
  • CleanVul - A Large-Scale, High-Quality Vulnerability Dataset: We introduced CleanVul, a new high-quality dataset derived from the application of the heuristic approach, containing 11,632 functions categorized as vulnerable and benign. We also showed that CleanVul maintains Correctness level (percentage of genuine vulnerable functions) comparable to existing high-quality datasets like SVEN and PrimeVul, enhancing the scale and reliability of data available for vulnerability detection research.
@@ 4. CleanVul: Dataset Curation
To create a high-quality vulnerability dataset using VulSifter, we conducted a comprehensive crawl of 127,063 repositories, resulting in the acquisition of 5,352,105 commits. Using the keyword-based approach proposed by Bui et al. (Bui et al., 2024), we identified 59,687 Vulnerability-Fix Commits (VFCs) from these repositories. We analyzed these VFCs using VulSifter, which processed code changes across multiple programming languages: 41,777 Java, 6,892 Python, 6,379 C, 4,180 JavaScript, 333 C#, and 126 C++ changes.
VulSifter assigns scores from 0 to 4 to each code change, where 0 indicates no relation to vulnerability fixes and 4 signifies a strong focus on fixing vulnerabilities. Table 2 presents the distribution of vulnerable functions across these threshold levels in our dataset CleanVul, evaluated both with () and without () heuristics. At threshold level 1, we identified 36,543 vulnerable functions with heuristics and 41,327 without. The number of detected functions decreases as the threshold increases: level 2 found 23,070 () and 25,789 () functions; level 3 identified 11,632 () and 12,847 () functions; and level 4 contained 8,337 () and 9,235 () functions.
@@ 6.1. RQ1: What is the correctness improvement of CleanVul compared to the uncleaned dataset?
With a threshold of 3, we obtain a dataset of 11,632 vulnerability-fixing changes, achieving a Correctness of 90.6% on the testing sample. Increasing the threshold to 4 results in 8,337 vulnerability-fixing changes with a perfect Correctness of 97.3%. These results match the Correctness levels of existing high-quality datasets SVEN (He and Vechev, 2023) (94.0%) and PrimeVul (Ding et al., 2024) (86.0%), while providing more samples compared to SVEN (803) and PrimeVul (6,968).
@@
Our CleanVul features a vulnerability score ranging from 0 to 4. To test the highest quality dataset we could obtain, we selected a threshold of 4. As mentioned in Section 6.1, this process resulted in a collection of 8,337 vulnerable and benign function pairs. To facilitate comparison with two other high-quality datasets, SVEN (He and Vechev, 2023) and PrimeVul (Ding et al., 2024), which exhibit high Correctness levels exceeding 85%, we also employed a balanced dataset setting for training and testing to eliminate other factors. It is important to note that PrimeVul contains 6,968 pairs of code changes exclusively in C and C++ code, while SVEN consists of 803 pairs of code changes, with equal parts C and C++ code and Python code. We partitioned the dataset into training, testing, and validation sets using a 7:1.5:1.5 ratio and subsequently compared their cross-dataset performance. The evaluation metrics employed were accuracy, precision, recall, and F1-score, which are presented for different LLMs trained on one dataset and tested on others, as illustrated in Table 5.
@@ 9.1. Vulnerability Datasets
Our work aims to bridge the gap between dataset size and quality. CleanVul achieves a correctness rate of 90.6% (comparable to SVEN and PrimeVul) while maintaining a larger scale of 11,632 functions. Importantly, CleanVul is derived from GitHub data rather than relying solely on NVD entries, making it a valuable complement to existing high-quality datasets.
@@ 10. Conclusion and Future Work
We created CleanVul, a new high-quality vulnerability dataset, by analyzing 5,352,105 commits from 127,063 GitHub repositories. Using VulSifter, we filtered out noise from vulnerability-fixing commits (VFCs) in this corpus. VulSifter has a configurable threshold that can be set based on data cleanliness requirements, allowing us to balance dataset size and quality. With a threshold of 3, CleanVul contains 11,632 vulnerable function pairs achieving 90.6% Correctness, while a stricter threshold of 4 yields 8,337 function pairs with 97.3% Correctness in our test sample. These results are comparable to established datasets such as SVEN (94.0% Correctness) and PrimeVul (86.0% Correctness), while overcoming their limitations. Unlike SVEN, which relies on manual analysis and contains only 803 samples, our approach is scalable. Additionally, while PrimeVul requires NVD entry links, our VulSifter can analyze all VFCs, including those without such links. Our evaluation of various LLMs fine-tuned on CleanVul revealed its superior generalization capabilities across different datasets. Notably, models fine-tuned on CleanVul significantly outperformed PrimeVul-trained models when tested on SVEN, achieving 64.87% accuracy compared to 55.75%. When evaluated on PrimeVul, our models demonstrated better performance than those trained on SVEN (58.09% vs 55.20%). Most remarkably, models trained on CleanVul and tested on PrimeVul achieved higher accuracy than models trained and tested on PrimeVul itself (58.09% vs 57.19%). These results underscore CleanVul’s effectiveness and robustness for model fine-tuning, particularly in its ability to capture generalizable vulnerability patterns that transfer well across different contexts.

### v6

# CleanVul: Toward High-Quality Function-Level Vulnerability Datasets via LLM-Based Noise Reduction

@@ Abstract.
To address these challenges, this paper introduces the first methodology that leverages the Large Language Model (LLM) with a heuristic enhancement to automatically identify vulnerability-fixing changes from VFCs, achieving an F1-score of 0.82. VulSifter was applied to a large-scale study, where we conducted a comprehensive crawl of 127,063 repositories on GitHub, resulting in the acquisition of 5,352,105 commits. The LLM with heuristic enhancement approach involves utilizing an LLM to comprehend code semantics and contextual information, while applying heuristics to filter out unrelated changes. We developed CleanVul, a high-quality dataset comprising 8,203 functions using our LLM heuristic enhancement approach, demonstrating Correctness (90.6%) comparable to established datasets such as SVEN (94.0%) (He and Vechev, 2023) and PrimeVul (86.0%) (Ding et al., 2024). In this context, Correctness is defined as the percentage of genuine vulnerable functions in the vulnerability dataset, which helps assess the effectiveness of VulSifter in identifying and filtering vulnerability-fixing changes. To evaluate the effectiveness of our CleanVul dataset, we conducted experiments focusing on fine-tuning various LLMs on the CleanVul dataset and other high-quality datasets. Our evaluation was conducted on multiple popular programming languages, including Java, Python, JavaScript, C#, C, and C++. Evaluation results reveal that LLMs fine-tuned on CleanVul not only exhibit enhanced accuracy but also superior generalization capabilities compared to those trained on uncleaned datasets. Specifically, models trained on CleanVul and tested on PrimeVul achieve accuracy higher than those trained and tested exclusively on PrimeVul, validating the effectiveness of our methodology.
@@ 1. Introduction
This paper introduces the first methodology to address the gap in accurately identifying vulnerability-related changes within VFCs. We first examine the reasons why many changes cataloged as vulnerability fixes do not pertain to actual security vulnerabilities. Our empirical study reveals that approximately 80% of non-vulnerability changes consist of test-related modifications and general bug fixes, with additional changes spanning support updates, code refactoring, and documentation updates. Building on this analysis, we developed VulSifter, a novel approach that combines LLM with heuristic filtering to automatically analyze VFCs and eliminate noise, resulting in a refined vulnerability dataset called CleanVul. By applying VulSifter, the Correctness of genuine vulnerability fixes in the dataset improved from 28.7% to 90.6%. Correctness is defined as the percentage of genuine vulnerable functions in the vulnerability dataset, and it helps assess the effectiveness of VulSifter in identifying and filtering vulnerability-fixing changes. CleanVul comprises 8,203 functions, both vulnerable and benign, and achieves a level of Correctness (90.6%) comparable to established datasets such as SVEN (He and Vechev, 2023) and PrimeVul (Ding et al., 2024). For comparison, the manually curated SVEN dataset contains 803 functions with 94.0% Correctness, while PrimeVul, which identifies vulnerabilities by matching function names with NVD descriptions, contains 6,968 NVD-sourced functions with 86.0% Correctness. As the first automated approach to improve dataset Correctness without requiring NVD linkage or other constraints, CleanVul represents a significant advancement in function-level vulnerability detection. It complements PrimeVul’s NVD-based approach by automatically filtering noise from GitHub VFCs to obtain high-quality vulnerability datasets.
@@
  • CleanVul - A Large-Scale, High-Quality Vulnerability Dataset: We introduced CleanVul, a new high-quality dataset derived from the application of the heuristic approach, containing 8,203 functions categorized as vulnerable and benign. We also showed that CleanVul maintains Correctness level (percentage of genuine vulnerable functions) comparable to existing high-quality datasets like SVEN and PrimeVul, enhancing the scale and reliability of data available for vulnerability detection research.
@@ 4. CleanVul: Dataset Curation
To create a high-quality vulnerability dataset using VulSifter, we conducted a comprehensive crawl of 127,063 repositories, resulting in the acquisition of 5,352,105 commits. Using the keyword-based approach proposed by Bui et al. (Bui et al., 2024), we identified 43,029 function changes from Vulnerability-Fix Commits (VFCs) from these repositories. We analyzed these VFCs using VulSifter, which processed code changes across multiple programming languages: 26,423 Java, 6,591 Python, 5,578 C, 4,000 JavaScript, 312 C#, and 125 C++ changes.
VulSifter assigns scores from 0 to 4 to each code change, where 0 indicates no relation to vulnerability fixes and 4 signifies a strong focus on fixing vulnerabilities. Table 2 presents the distribution of vulnerable functions across these threshold levels in our dataset CleanVul, evaluated both with () and without () heuristics. At threshold level 1, we identified 26,549 vulnerable functions with heuristics and 29,841 without. The number of detected functions decreases as the threshold increases: level 2 found 18,465 () and 16,287 () functions; level 3 identified 8,203 () and 9,031 () functions; and level 4 contained 6,371 () and 7,023 () functions.
@@ 6.1. RQ1: What is the correctness improvement of CleanVul compared to the uncleaned dataset?
With a threshold of 3, we obtain a dataset of 8,203 vulnerability-fixing changes, achieving a Correctness of 90.6% on the testing sample. Increasing the threshold to 4 results in 6,371 vulnerability-fixing changes with a perfect Correctness of 97.3%. These results match the Correctness levels of existing high-quality datasets SVEN (He and Vechev, 2023) (94.0%) and PrimeVul (Ding et al., 2024) (86.0%), while providing more samples compared to SVEN (803) and PrimeVul (6,968).
@@
Our CleanVul features a vulnerability score ranging from 0 to 4. To test the highest quality dataset we could obtain, we selected a threshold of 4. As mentioned in Section 6.1, this process resulted in a collection of 6,371 vulnerable and benign function pairs. To facilitate comparison with two other high-quality datasets, SVEN (He and Vechev, 2023) and PrimeVul (Ding et al., 2024), which exhibit high Correctness levels exceeding 85%, we also employed a balanced dataset setting for training and testing to eliminate other factors. It is important to note that PrimeVul contains 6,968 pairs of code changes exclusively in C and C++ code, while SVEN consists of 803 pairs of code changes, with equal parts C and C++ code and Python code. We partitioned the dataset into training, testing, and validation sets using a 7:1.5:1.5 ratio and subsequently compared their cross-dataset performance. The evaluation metrics employed were accuracy, precision, recall, and F1-score, which are presented for different LLMs trained on one dataset and tested on others, as illustrated in Table 5.
@@ 9.1. Vulnerability Datasets
Our work aims to bridge the gap between dataset size and quality. CleanVul achieves a correctness rate of 90.6% (comparable to SVEN and PrimeVul) while maintaining a larger scale of 8,203 functions. Importantly, CleanVul is derived from GitHub data rather than relying solely on NVD entries, making it a valuable complement to existing high-quality datasets.
@@ 10. Conclusion and Future Work
We created CleanVul, a new high-quality vulnerability dataset, by analyzing 5,352,105 commits from 127,063 GitHub repositories. Using VulSifter, we filtered out noise from vulnerability-fixing commits (VFCs) in this corpus. VulSifter has a configurable threshold that can be set based on data cleanliness requirements, allowing us to balance dataset size and quality. With a threshold of 3, CleanVul contains 8,203 vulnerable function pairs achieving 90.6% Correctness, while a stricter threshold of 4 yields 6,371 function pairs with 97.3% Correctness in our test sample. These results are comparable to established datasets such as SVEN (94.0% Correctness) and PrimeVul (86.0% Correctness), while overcoming their limitations. Unlike SVEN, which relies on manual analysis and contains only 803 samples, our approach is scalable. Additionally, while PrimeVul requires NVD entry links, our VulSifter can analyze all VFCs, including those without such links. Our evaluation of various LLMs fine-tuned on CleanVul revealed its superior generalization capabilities across different datasets. Notably, models fine-tuned on CleanVul significantly outperformed PrimeVul-trained models when tested on SVEN, achieving 64.87% accuracy compared to 55.75%. When evaluated on PrimeVul, our models demonstrated better performance than those trained on SVEN (58.09% vs 55.20%). Most remarkably, models trained on CleanVul and tested on PrimeVul achieved higher accuracy than models trained and tested on PrimeVul itself (58.09% vs 57.19%). These results underscore CleanVul’s effectiveness and robustness for model fine-tuning, particularly in its ability to capture generalizable vulnerability patterns that transfer well across different contexts.